OpenStudy (anonymous):

can anyone answer me..the formations of definite integral of f(x) on [a,b] ?

OpenStudy (anonymous):

insufficient information

OpenStudy (anonymous):

i guess: let f'(x) = integral of f(x) ==> f'(x)(b-a)

OpenStudy (anonymous):

owh...ok thanks

OpenStudy (anonymous):

f'(x) is the derivative, not the integral

OpenStudy (anonymous):

not if u let it = intergral :P

OpenStudy (anonymous):

i just couldnt think of a way to represent it. i guess i should have said let g(x) = integral f(x)

OpenStudy (anonymous):

is that the formation of definite integral? i really dont understand in...:(

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

tbh, the question seems way to general to be a question. Is that the exact wording?

OpenStudy (anonymous):

the question : discuss (in detail) the formations of definite integral of f(x) on [a,b].

OpenStudy (across):

Is this what you're looking for?\[\int_{a}^{b}f(x)dx=\lim_{n\to\infty}\sum_{i=1}^{n}f(c_{i})\Delta x_{i}\]

OpenStudy (anonymous):

ya..but how to explain it?

OpenStudy (across):

Where\[\Delta x_i=\frac{b-a}{n},\]\[c_i=a+\left(\frac{b-a}{n}\right)i.\]You could use something else for the second equation above.

OpenStudy (anonymous):

owh..thanks! you help me a lot ^^

OpenStudy (across):

You could say that the integral of a function within a range (a, b) is defined to be the summation of infinitely-many rectangles with (often) equally sized bases, which are partitions of the range (a, b), (having infinitely small length) and height defined as the function value evaluated at each partition.

OpenStudy (anonymous):

perfect! thanks